HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

NON-LTE SPECTROSCOPY OF METHANE RECORDED IN A HYPERSONIC FLOW

E. DUDÁS, N. SUAS-DAVID, L. ROLLAND, \underline{R. GEORGES} , IPR, Université de Rennes, UMR6251 CNRS, F-35042, France; B. VISPOEL, Research Unit Lasers

and Spectroscopies (LLS), Institute of Life, Earth and Environment (ILEE), University of Namur (UNamur), 61 rue de Bruxelles, B-5000, Namur, Belgium; R. R. GAMACHE, Department of Environmental, Earth, and Atmospheric Sciences University of Massachusetts Lowell, Lowell, MA 01854 USA; M. REY, GSMA, UMR 7331, Université de Reims, Reims, France; V. TYUTEREV, A. NIKITIN, V.E. Zuev Institute of Atmospheric Optics SB RAS, Tomsk, 634055, Russia; S. KASSI, LIPhy, UMR 5588, Université Grenoble Alpes, Saint Martin d’Hères,

France

Laboratory spectroscopic data is essential for the modeling of hot exoplanet atmospheres, since molecules such as methane, a major component of hot-Jupiter-type exoplanet atmospheres, have a complex vibrational energy structure that makes computational predictions difficult at high temperatures for ro-vibrational transitions involving highly excited vibrational sates. To better inform line lists used in radiative transfer modelling, the ro-vibrational spectrum of methane has been recorded in the tetradecad region between 1.7 and 1.65 μm (5880-6060 cm^{−1}) through non-local thermodynamic equilibrium (non-LTE) cavity ringdown spectroscopy (CRDS)^{1}. Non-LTE conditions, characterized by a low rotational temperature (39 K) and a high vibrational temperature (up to 1130 K), have been obtained by hypersonic expansion of a pre-heated mixture of argon and methane in a contoured Laval nozzle. The high vibrational temperature increases the intensity of new hot bands, while the very low rotational temperature greatly simplifies their rotational structure, thus facilitating their identification. A close comparison of the recorded CRDS data to the TheoReTS database reveals both inefficient vibrational relaxation between polyads and efficient vibrational relaxation between vibrational states forming a polyad. These effects result in an overpopulation of the lowest vibrational energy level of each polyad, an effect not widely currently incorporated in non-LTE radiative transfer models. A series of new hot band transitions originating from the pentad and octad polyads were assigned and are provided as a line list for use in future databases^{2}.

1doi:10.1063/5.0003886, E. Dudás et al., J. Chem. Phys., 152, 134201 (2020).

2doi:10.1016/j.icarus.2022.115421, E. Dudás et al., Icarus, 394, 115421 (2023).

p-number: p142  Submitted on Thu, 15 Jun 2023 11:31:14 +0200